# On the Convergence Analysis of Muon

### v1

@@ Abstract
The majority of parameters in neural networks are naturally represented as matrices. However, most commonly used optimizers treat these matrix parameters as flattened vectors during optimization, potentially overlooking their inherent structural properties. Recently, an optimizer called Muon has been proposed, specifically designed to optimize matrix-structured parameters. Extensive empirical evidence shows that Muon can significantly outperform traditional optimizers when training neural networks. Nonetheless, the theoretical understanding of Muon’s convergence behavior and the reasons behind its superior performance remain limited. In this work, we present a comprehensive convergence rate analysis of Muon and its comparison with Gradient Descent (GD). We further characterize the conditions under which Muon can outperform GD. Our theoretical results reveal that Muon can benefit from the low-rank and approximate blockwise diagonal structure of Hessian matrices – phenomena widely observed in practical neural network training. Our experimental results support and corroborate the theoretical findings.
@@ 1 Introduction
Modern neural networks – such as large language models (LLMs) Adler et al. (2024) – typically consist of a massive amount of parameters and require significant computational resources for training. As a result, designing optimizers that can efficiently train such models has become an important and valuable research problem. Note that most of the parameters in neural networks are naturally represented as matrices – for example, weight matrices in fully connected layers, or the query, key, and value matrices in attention mechanisms. However, most widely adopted optimization algorithms such as Stochastic Gradient Descent (SGD), Adam (Diederik, 2014), and their variants often treat these matrices as flattened vectors, potentially overlooking their inherent structural properties.
@@
Recently, Muon, an optimizer specifically designed for matrix-structured parameters, was proposed in Jordan et al. (2024). The key idea is to update the matrix parameters along an orthogonalized version of their gradients at each iteration (Algorithm 1). Empirical results across various studies and network scales consistently demonstrate the superior performance of Muon in optimizing neural networks with matrix parameters (Jordan et al., 2024; Liu et al., 2025a; An et al., 2025; Liu et al., 2025b; Ahn and Xu, 2025). However, the convergence behavior of Muon and the underlying mechanisms that contribute to its advantage over traditional gradient-descent-based optimizers, such as Gradient Descent (GD) and SGD, are not yet fully understood. Therefore, in this work, we aim to bridge this gap by providing a comprehensive theoretical analysis of Muon’s convergence properties, along with detailed comparisons to GD-based algorithms.
The stepsize for GD usually depends on the Lipschitz smoothness constant (the max singular value of Hessians), which can change rapidly or be very large. This makes it difficult to tune the stepsize. Instead of the max singular value of Hessians (Frobenius norm Lipschitz constant (3.1)), by analyzing the convergence behaviors of Muon without assuming the uniform Lipschitz smoothness, we show that the behavior of Muon depends on the average of the global information of the Hessian matrices during training, which can be viewed as smoothing over steps and singular values compared with the Lipschitz constant of GD (the max singular value of Hessians), which gives the potential advantages of Muon over GD.
Our main contributions can be summarized as follows:
  • In the nonconvex setting, we rigorously establish the convergence guarantees for Muon in both deterministic and stochastic cases. We first provide detailed comparisons with GD, demonstrating that Muon can benefit from the low-rank (Sagun et al., 2016, 2017; Wu et al., 2020; Papyan, 2020; Yao et al., 2020) and approximately blockwise diagonal structure of Hessian matrices (Dong et al., 2025; Zhang et al., 2024b, a; Collobert, 2004), which are widely observed phenomena in practical neural network training. Moreover, we analyze the convergence of Muon without assuming the uniform Lipschitz smoothness (Theorem 4.8). We show that the convergence rate of Muon can be related to the average of the global information of the Hessian matrices during training. We characterize the conditions under which Muon outperforms GD, and validate these theoretical findings through experiments on neural network training.
  • To further study the advantages of Muon compared to GD, we consider star convex functions (4.9), for which we can study the convergence rate of function value. In the star convex setting, under mild assumptions, we show that Muon can achieve a convergence rate comparable to GD’s. Furthermore, we prove that when the Hessian matrices exhibit relatively low-rank and approximately blockwise diagonal structures, Muon can potentially outperform GD. Our experiments on quadratic functions validate and support these theoretical findings.
@@ 2 Related Work
Muon was originally proposed in Jordan et al. (2024). (Jordan et al., 2024) also demonstrated the superior performance of Muon on various models and datasets. Subsequently, Liu et al. (2025a) improved the Muon by incorporating techniques such as weight decay and adjusting the per-parameter update scale, and showed that Muon can outperform Adam when optimizing large language models (LLMs). One possible explanation for Muon’s superior performance is provided by Bernstein and Newhouse (2024), whose work shows that Muon’s update direction corresponds to the steepest descent under the spectral norm constraint. However, during the writing of this paper, we note that there are some concurrent works (Li and Hong, 2025; An et al., 2025; Kovalev, 2025) that also analyze the convergence behavior of Muon. For example, Li and Hong (2025) analyze the Frobenius norm convergence (Definition 3.4) of Muon under Frobenius norm Lipschitz smooth (3.1). An et al. (2025) analyze the convergence of the Simplified Muon (Algorithm 2) with 4.5. Kovalev (2025) provide the convergence analysis of Muon with 3.2. However, compared to these prior works (Li and Hong, 2025; An et al., 2025; Kovalev, 2025), our work provides a more comprehensive theoretical analysis, i.e. we analyze the convergence of Muon under various smoothness assumptions (3.1, 3.2, 4.5), including scenarios without uniform Lipschitz smoothness (4.7). Additionally, we offer detailed comparisons with GD and characterize the conditions under which Muon outperforms GD, and validate these conditions through experiments. Therefore, our work presents distinct contributions and insights. We believe that all these works collectively contribute to a deeper understanding of Muon.
@@
Low-rank and approximate blockwise diagonal structure of Hessian. Extensive prior works (Collobert, 2004; Zhang et al., 2024b, a; Dong et al., 2025) have shown that the Hessian of a neural network tends to exhibit a blockwise diagonal structure, with each block corresponding to an individual neuron, both theoretically and empirically. Recently, Zhang et al. (2024b, a) numerically confirmed this property in small Transformers, and Dong et al. (2025) provided theoretical explanations for why such a blockwise diagonal structure emerges in neural network Hessians. In addition, numerous studies (Sagun et al., 2016, 2017; Wu et al., 2020; Papyan, 2020; Yao et al., 2020) have also observed that neural network Hessians are typically low-rank, i.e., their effective ranks, which can be measured by the ratio $\|H\|_{*}/\|H\|_{{\rm op}}$, can be significantly smaller than their dimensionality.
@@
Other structured optimization methods. Although commonly used optimizers for training deep neural networks, such as SGD, Adam (Diederik, 2014), and AdamW (Loshchilov and Hutter, 2017), typically treat structured parameters (e.g., matrices) as flattened vectors, however, in recent years, there has been growing interest in designing structured optimizers that explicitly leverage the inherent structure of parameters. For instance, Adafactor (Shazeer and Stern, 2018), LAMB (You et al., 2019), and Adam-mini (Zhang et al., 2024b) incorporate matrix- or layer-level structure to reduce memory consumption. KFAC (Martens and Grosse, 2015) and TNT (Ren and Goldfarb, 2021) approximate the Fisher matrix to implement natural gradient methods. Shampoo (Gupta et al., 2018) is specifically designed for matrix or tensor parameters and can be viewed as an approximation to the full-matrix preconditioner in AdaGrad (Duchi et al., 2011). Morwani et al. (2024) provided additional theoretical analyses and modifications to Shampoo. Recently, SOAP (Vyas et al., 2024) was introduced, which combines the ideas of Adam and Shampoo. Galore (Zhao et al., 2024) was proposed to exploit the low-rank structure of gradients for memory efficiency. Additionally, Liu et al. (2025b) proposed COSMOS, which can be seen as a combination of SOAP and Muon. More recently, Xie et al. (2025) and An et al. (2025) introduced ASGO (One-Sided Shampoo), which only applies a one-sided preconditioner within the Shampoo. Xie et al. (2025); An et al. (2025) showed that ASGO(One-Sided Shampoo) can achieve better convergence rate than the original Shampoo.
@@ 3 Preliminaries
Notation. For a vector $v$, we denote its $l_{2}$ norm as $\|v\|_{2}$. For a matrix $A\in\mathbb{R}^{m\times n}$, we denote its nuclear norm as $\|A\|_{*}$, spectral norm as $\|A\|_{{\rm op}}$, Frobenius norm as $\|A\|_{\rm F}$, and $\Lambda$ norm as $\|A\|_{\Lambda}=\sqrt{\text{tr}\left(A\Lambda A^{\top}\right)}$ where $\Lambda\in\mathbb{R}^{n\times n}$ is a real symmetric positive definite matrix. For $A\in\mathbb{R}^{m\times n}$ with rows $a_{1},\ldots,a_{m}$, we define $\mathsf{vec}(A)=(a_{1}\,a_{2}\,\cdots\,a_{m})^{\top}\in\mathbb{R}^{mn}$ as the vectorization of $A$. We denote $I_{d}$ as the identity matrix with $d$ dimension. For matrix $A\in\mathbb{R}^{m\times n}$, $B\in\mathbb{R}^{m^{\prime}\times n^{\prime}}$, we denote their Kronecker product as $A\otimes B\in\mathbb{R}^{mm^{\prime}\times nn^{\prime}}$. We define the operator norm of a third‑order tensor $H\in\mathbb{R}^{d_{1}\times d_{2}\times d_{3}}$ as $\|H\|_{{\rm op}}=\max_{\|u\|_{2}=\|v\|_{2}=\|w\|_{2}=1}|\sum_{i=1}^{d_{1}}\sum_{j=1}^{d_{2}}\sum_{k=1}^{d_{3}}H_{ijk}u_{i}v_{j}w_{k}|$, where $u\in\mathbb{R}^{d_{1}},v\in\mathbb{R}^{d_{2}},w\in\mathbb{R}^{d_{3}}$.
@@
We denote $f^{*}=\inf_{W}f(W)$, $r=\min\{m,n\},f_{v}(w)=f(W)$, and $w=\mathsf{vec}(W)\in\mathbb{R}^{mn}$. We assume $f^{*}>-\infty$. We consider the following smoothness assumptions in this paper.
@@
The first is the standard Frobenius norm-based smoothness 3.1, which serves as a natural extension of the conventional $l_{2}$-norm smoothness for functions with vector parameters to functions with matrix parameters. The second is a spectral norm smoothness 3.2, which accounts for the distinct structure of matrix parameters. In this work, we establish the convergence of Muon under both assumptions.
@@
For a matrix $A$ with rank $r$, there is a well-known relationship between its Frobenius norm and nuclear norm: $\|A\|_{\rm F}\leq\|A\|_{*}\leq\sqrt{r}\|A\|_{\rm F}$. Thus, we have following proposition.
@@ 3.1 Muon
In this subsection, we introduce the Muon algorithm. The original idea of Muon in Jordan et al. (2024) is to orthogonalize the update matrix. For example, suppose the original update direction is $G_{t}$ with rank $r_{t}$ and the singular value decomposition (SVD) of $G_{t}$ is $U_{t}S_{t}V_{t}^{\top}$, where $S_{t}\in\mathbb{R}^{r_{t}\times r_{t}}$ is the diagonal matrix of the singular values of $G_{t}$, $U_{t}\in\mathbb{R}^{m\times r_{t}}$ and $V_{t}\in\mathbb{R}^{n\times r_{t}}$ are the left and right singular vector matrices of $G_{t}$, respectively. Then, in Muon, the update matrix will be $U_{t}V_{t}^{\top}$, which is the nearest semi-orthogonal matrix to the original $G_{t}$. Following this main idea, we have Algorithm 2, which can be viewed as the simplest form of Muon.
Additionally, if we add momentum and conduct orthogonalization over the momentum direction, we have Algorithm 1, which is the main algorithm we will analyze in the stochastic setting. In fact, our later experimental and theoretical results show that Algorithm 1 and 2 have very similar performance in the deterministic setting, while additional momentum can be helpful in the stochastic setting (See Figure 1).
Furthermore, in practice, computing the SVD is very costly. Therefore, a common practical implementation of Muon uses Newton–Schulz iterations Bernstein and Newhouse (2024); Jordan et al. (2024) to approximate the orthogonalization process. An introduction of the Muon with Newton–Schulz iterations can be found in Appendix E. However, our later experiments show that the performance of the SVD-based version of Muon is similar to the Newton–Schulz-based version, with the primary difference being the higher computational cost of the SVD procedure. (See Figure 1). Therefore, without loss of generality, we primarily analyze the convergence properties of Algorithm 1 and 2 using SVD in this paper.
(a) Deterministic
(b) Stochastic
Figure 1: Comparison of (S)GD, Adam, Muon (using Newton–Schulz iterations with momentum), Muon_withoutM (using Newton–Schulz iterations without momentum), and Muon_SVD (Algorithm 1). In the deterministic setting (a), loss is defined and trained over a fixed subset of CIFAR-10 (Krizhevsky et al., 2009). In the stochastic setting (b), training utilizes mini-batches randomly sampled from the complete CIFAR-10 training set. The loss is evaluated on the entire CIFAR-10 training set.
4 Convergence of Muon
4.1 Nonconvex case
In this subsection, we analyze the convergence guarantees of Muon in the nonconvex setting with various smoothness assumptions. All the proofs of theorems in this subsection can be found in Appendix B. We first prove the convergence of Muon with the Frobenius norm Lipschitz smoothness assumption.

### v2

@@ Abstract
The majority of parameters in neural networks are naturally represented as matrices. However, most commonly used optimizers treat these matrix parameters as flattened vectors during optimization, potentially overlooking their inherent structural properties. Recently, an optimizer called Muon has been proposed, specifically designed to optimize matrix-structured parameters. Extensive empirical evidence shows that Muon can significantly outperform traditional optimizers when training neural networks. Nonetheless, the theoretical understanding of Muon’s convergence behavior and the reasons behind its superior performance remain limited. In this work, we present a comprehensive convergence rate analysis of Muon and its comparison with Gradient Descent (GD). We characterize the conditions under which Muon can outperform GD. Our theoretical results reveal that Muon can benefit from the low-rank structure of Hessian matrices, a phenomenon widely observed in practical neural network training. Our experimental results support and corroborate the theoretical findings.
@@ 1 Introduction
Modern neural networks, such as large language models (LLMs) (Adler et al., 2024), typically consist of a massive amount of parameters and require significant computational resources for training. As a result, designing optimizers that can efficiently train such models has become an important and valuable research problem. Note that most of the parameters in neural networks are naturally represented as matrices, for example, weight matrices in fully connected layers, or the query, key, and value matrices in attention mechanisms. However, most widely adopted optimization algorithms such as Stochastic Gradient Descent (SGD), Adam (Kingma and Ba, 2015), and their variants often treat these matrices as flattened vectors, potentially overlooking their inherent structural properties.
@@
Recently, Muon, an optimizer specifically designed for matrix-structured parameters, was proposed in Jordan et al. (2024). The key idea is to update the matrix parameters along an orthogonalized version of their gradients at each iteration (Algorithm 1). Empirical results across various studies and network scales consistently demonstrate the superior performance of Muon in optimizing neural networks with matrix parameters (Jordan et al., 2024; Liu et al., 2025a; An et al., 2025; Liu et al., 2025b; Ahn and Xu, 2025). However, the convergence behavior of Muon and the underlying mechanisms that contribute to its advantage over traditional gradient descent-based optimizers, such as GD and SGD, are not yet fully understood. Therefore, in this work, we aim to bridge this gap by providing a comprehensive theoretical analysis of Muon’s convergence properties, along with detailed comparisons to GD-based algorithms.
In classical optimization theory, the stepsize of GD typically depends on the Lipschitz smoothness constant $L$ of the Frobenius norm, namely the maximum singular value of the Hessian. However, this quantity can be extremely large or vary rapidly during neural network training, making the stepsize of GD difficult to tune in practice. In this work, we analyze the convergence behavior of Muon not only under the standard Lipschitz smoothness constant $L$, but also under the spectral norm Lipschitz smoothness constant $L_{*}$, and even without assuming uniform Lipschitz smoothness. We show that the convergence of Muon can depend on $J$, which can be viewed as an average of the global Hessian information throughout training. Compared with $L$, which is determined by the maximum singular value of the Hessian, $J$ can be interpreted as averaging across both iterations and singular values. Such a smoothing effect provides potential advantages for Muon over GD-based algorithms, especially when the Hessians exhibits low-rank structures, a phenomenon widely observed in practical neural network training. Our experimental results on quadratic regressions and neural network trainings validate these theoretical findings.
@@ 2 Related Work
Muon was originally proposed in Jordan et al. (2024), which demonstrated the superior performance of Muon on various models and datasets. Subsequently, Liu et al. (2025a) improved the Muon by incorporating techniques such as weight decay and adjusting the per-parameter update scale, and showed that Muon can outperform Adam when optimizing large language models (LLMs). One possible explanation for Muon’s superior performance is provided by Bernstein and Newhouse (2024), which showed that Muon’s update direction corresponds to the steepest descent under the spectral norm constraint. Moreover, we note that there are some recent works (Li and Hong, 2025; Pethick et al., 2025; An et al., 2025; Kovalev, 2025) that also analyze the convergence behavior of Muon. For example, Li and Hong (2025) analyzed the Frobenius norm convergence (Definition 3.4) of Muon under Frobenius norm Lipschitz smooth (3.1). An et al. (2025) analyzed the convergence of the Simplified Muon (Algorithm 2) with 4.11. Pethick et al. (2025) viewed Muon as one of the linear minimization oracle (lmo) based family algorithms and Kovalev (2025) viewed Muon as a special case of first-order trust-region optimization methods. Both Pethick et al. (2025); Kovalev (2025) provided the convergence analysis of Muon with 3.2.
Recently, several works have also investigated the superiority of Muon from theoretical perspectives. Wang et al. (2025) demonstrated that Muon can outperform Adam in tail-end associative memory learning, showing that spectral orthogonalization enables more balanced updates across singular modes. Su (2025) introduced an isotropic curvature model to analyze matrix-structured optimization updates. Davis and Drusvyatskiy (2025) analyzed the squared nuclear-to-Frobenius ratio of the gradient relative to the stable rank of the incoming activations to characterize when Muon outperforms GD, and show that this condition holds in random-feature models, deep networks, and transformer training. Ma et al. (2026) analyzed matrix factorization and in-context learning of linear transformers, showing that for both problems Muon achieves iteration complexity independent of the condition number, thereby outperforming GD and Adam.
Compared to these prior works, our work provides a more comprehensive convergence analysis of Muon under various smoothness assumptions (3.1, 3.2, 4.11), including scenarios without uniform Lipschitz smoothness (4.13). Additionally, we offer detailed comparisons with GD and characterize the conditions under which Muon outperforms GD, and validate these conditions through experiments. Therefore, our work presents distinct contributions and insights. We believe that all these works collectively contribute to a deeper understanding of Muon.
@@
Low-rank and approximate blockwise diagonal structure of Hessian. Extensive prior works (Collobert, 2004; Zhang et al., 2024b, a; Dong et al., 2025) have shown, both theoretically and empirically, that the Hessian of a neural network tends to exhibit a blockwise diagonal structure with each block corresponding to an individual neuron. Recently, Zhang et al. (2024b, a) numerically confirmed this property in small Transformers, and Dong et al. (2025) provided theoretical explanations for why such a blockwise diagonal structure emerges in neural network Hessians. In addition, many studies (Sagun et al., 2016, 2017; Wu et al., 2020; Papyan, 2020; Yao et al., 2020) have also observed that neural network Hessians are typically low-rank, i.e., their effective ranks, which can be measured by the ratio $\|H\|_{*}/\|H\|_{{\rm op}}$, can be significantly smaller than their dimensionality.
@@
Additional Structured Optimization Methods and Muon Variants. Although optimizers commonly used for training deep neural networks, such as SGD and Adam, typically treat structured parameters (e.g., matrices) as flattened vectors, in recent years, there has been growing interest in designing structured optimizers that explicitly leverage the inherent structure of parameters. For instance, Adafactor (Shazeer and Stern, 2018), LAMB (You et al., 2019), and Adam-mini (Zhang et al., 2024b) incorporate matrix- or layer-level structure to reduce memory consumption. KFAC (Martens and Grosse, 2015) and TNT (Ren and Goldfarb, 2021) approximate the Fisher matrix to implement natural gradient methods. Shampoo (Gupta et al., 2018) is specifically designed for matrix or tensor parameters and can be viewed as an approximation to the full-matrix preconditioner in AdaGrad (Duchi et al., 2011). Morwani et al. (2024) provided additional theoretical analyses and modifications to Shampoo. Recently, SOAP (Vyas et al., 2024) was introduced, which combines the ideas of Adam and Shampoo. Galore (Zhao et al., 2024) was proposed to exploit the low-rank structure of gradients for memory efficiency. Additionally, Liu et al. (2025b) proposed COSMOS, which can be seen as a combination of SOAP and Muon. More recently, Xie et al. (2025) and An et al. (2025) introduced ASGO (One-Sided Shampoo), which only applies a one-sided preconditioner within the Shampoo. Xie et al. (2025); An et al. (2025) showed that ASGO can achieve better convergence rate than the original Shampoo. Crawshaw et al. (2025) provide a unified non-Euclidean steepest descent framework for neural network optimization, revealing Muon and its variants as instances of product-norm-based gradient methods and introducing more robust alternatives such as MuonMax-Momo. Liu et al. (2025c) and Qian et al. (2025) incorporate momentum variance reduction into Muon-type optimizers, achieving provably faster convergence rates and improved empirical performance.
@@ 3 Preliminaries
Notation. For a vector $v$, we denote its $l_{2}$ norm as $\|v\|_{2}$. For a matrix $A\in\mathbb{R}^{m\times n}$, we denote its nuclear norm as $\|A\|_{*}$, spectral norm as $\|A\|_{{\rm op}}$, Frobenius norm as $\|A\|_{\rm F}$, and $\Lambda$ norm as $\|A\|_{\Lambda}=\sqrt{\text{tr}\left(A\Lambda A^{\top}\right)}$ where $\Lambda\in\mathbb{R}^{n\times n}$ is a real symmetric positive definite matrix. For $A\in\mathbb{R}^{m\times n}$ with rows $a_{1},\ldots,a_{m}$, we define $\overline{\mathrm{vec}}(A)=(a_{1}\,a_{2}\,\cdots\,a_{m})^{\top}\in\mathbb{R}^{mn}$ as the vectorization of $A$. We denote $I_{d}$ as the identity matrix with dimension $d$. For matrix $A\in\mathbb{R}^{m\times n}$, $B\in\mathbb{R}^{m^{\prime}\times n^{\prime}}$, we denote their Kronecker product as $A\otimes B\in\mathbb{R}^{mm^{\prime}\times nn^{\prime}}$. We define the operator norm of a third‑order tensor $H\in\mathbb{R}^{d_{1}\times d_{2}\times d_{3}}$ as $\|H\|_{{\rm op}}=\max_{\|u\|_{2}=\|v\|_{2}=\|w\|_{2}=1}|\sum_{i=1}^{d_{1}}\sum_{j=1}^{d_{2}}\sum_{k=1}^{d_{3}}H_{ijk}u_{i}v_{j}w_{k}|$, where $u\in\mathbb{R}^{d_{1}},v\in\mathbb{R}^{d_{2}},w\in\mathbb{R}^{d_{3}}$.
3.1 Optimization problem
@@
We denote $f^{*}=\inf_{W}f(W)$, $r=\min\{m,n\},f_{v}(w)=f(W)$, and $w=\overline{\mathrm{vec}}(W)\in\mathbb{R}^{mn}$. We assume $f^{*}>$-$\infty$. For initial point $W_{0}$, we denote $\Delta=f(W_{0})-f^{*}$, and $D_{F}=\|W_{0}-W^{*}\|_{F}$. We consider the following smoothness assumptions in this paper.
@@
3.1 is the standard Frobenius norm-based smoothness, which serves as a natural extension of the conventional $l_{2}$-norm smoothness for functions with vector parameters to functions with matrix parameters. 3.2 is a spectral norm smoothness, which accounts for the distinct structure of matrix parameters. In this work, we establish the convergence of Muon under both assumptions.
@@
For a matrix $A$ with rank $r$, there is a well-known relationship between its Frobenius norm and nuclear norm: $\|A\|_{\rm F}\leq\|A\|_{*}\leq\sqrt{r}\|A\|_{\rm F}$. Thus, we have the following proposition.
@@ 3.2 Muon
In this subsection, we introduce the Muon algorithms. The original idea of Muon in Jordan et al. (2024) is to orthogonalize the update matrix. For example, suppose the original update direction is $G_{t}$ with rank $r_{t}$ and the singular value decomposition (SVD) of $G_{t}$ is $U_{t}S_{t}V_{t}^{\top}$, where $S_{t}\in\mathbb{R}^{r_{t}\times r_{t}}$ is the diagonal matrix of the singular values of $G_{t}$, $U_{t}\in\mathbb{R}^{m\times r_{t}}$ and $V_{t}\in\mathbb{R}^{n\times r_{t}}$ are the left and right singular vector matrices of $G_{t}$, respectively. Then, in Muon, the update matrix will be $U_{t}V_{t}^{\top}$, which is the nearest semi-orthogonal matrix to the original $G_{t}$. Following this main idea, we have Algorithm 2, which can be viewed as the simplest form of Muon. Additionally, if we add momentum and conduct orthogonalization over the momentum direction, we have Algorithm 1, which is the main algorithm we will analyze in the stochastic setting. In practice, computing the SVD is very costly. Therefore, a common practical implementation of Muon uses Newton–Schulz iterations (Bernstein and Newhouse, 2024; Jordan et al., 2024) to approximate the orthogonalization process. In this paper, we primarily analyze the theoretical properties of Algorithm 1 and 2 using SVD.
4 Convergence Analysis of Muon
In this section, we analyze the convergence behaviors of Muon under various smoothness assumptions and compare them with GD. The proofs of theorems in this section can be found in Appendix B and Appendix C.
4.1 Frobenius norm Lipschitz smooth
We first analyze the convergence of Muon with the conventional Frobenius norm Lipschitz smoothness assumption (3.1).